OpenStudy (anonymous):

PLEASE HELP! I DON'T UNDERSTAND! <1 and <2 are linear pair. m<1=x-30, and m<2=x+76. find the measure of each angle? 1) <1=67,<2=123 2)<1=37,<2=153 3)<1=37,<2=143 4)<1=67,<2=113

Directrix (directrix):

@karinahorses A linear pair of angles are two adjacent angles whose exterior sides form a straight angle. A straight angle has measure 180 degrees.

Directrix (directrix):

m<1 + m<2 = 180 x-30 + x+76 = 180 @karinahorses Solve the equation for x. Post what you get and we'll compare answers. Okay? Then, we have to find the angle measures.

OpenStudy (jdoe0001):



OpenStudy (anonymous):

I have a very hard time with this stuff? I've never gotten it would it be 4? (or d)

OpenStudy (jdoe0001):

well, tis understandable karinahorses

Directrix (directrix):

@karinahorses I intended to followup with you on this problem earlier today when the OpenStudy site went down. So, let us continue now. m<1 + m<2 = 180 x-30 + x+76 = 180 (Combine like terms on the left side of the equation) 2x + 46 = 180 (Subtract 46 from each side of the equation) 2x = 134 (Divide both sides by 2) x = 67 Now, find the measures of angles 1 and 2 m<1=x-30 = 67 - 30 = ? @katrinahorses and m<2=67+76 = ? Note: When you get the measure of the two angles, check to see if they add to 180.

OpenStudy (anonymous):

3)<1=37,<2=143 After a while I have found my answer THANK YOU SO MUCH!:)

OpenStudy (anonymous):

I knew it was either 3 or 4 (C or D) and I read it and looked in my book to find out more about this certain question.

Directrix (directrix):

Alrighty. You are welcome. The textbook is always a good first resource when working problems, especially Geometry problems.

OpenStudy (anonymous):

Lol I read it at first Just didn"t get it and being a quiz I was try not to resort to my book but I looked over all my answers and only got 1 other wrong with all 12 question and I totally guessed on that one:)